### v1

# Bracketing Transduction Grammar BPE Byte Pair Encoding

BTG
Bracketing Transduction Grammar
BPE
Byte Pair Encoding
BSL
British Sign Language
BLEU
Bilingual Evaluation Understudy
BOBSL
BBC-Oxford British Sign Language
BLSTM
Bidirectional Long Short-Term Memory
BSLCPT
BSL Corpus T
CNN
Convolutional Neural Network
CRF
Conditional Random Field
CSLR
Continuous Sign Language Recognition
CTC
Connectionist Temporal Classification
C4A
Content4All
DL
Deep Learning
DGS
German Sign Language - Deutsche Gebärdensprache
DSGS
Swiss German Sign Language - Deutschschweizer Gebärdensprache
DTW
Dynamic Time Warping
DTW-MJE
Dynamic Time Warping Mean Joint Error
FC
Fully Connected
FF
Feed Forward
fps
frames per second
GAN
Generative Adversarial Network
GPU
Graphics Processing Unit
GRU
Gated Recurrent Unit
G2PT
Gloss-to-Pose Transformer
G2P
Gloss-to-Pose
G2S
Gloss-to-Sign
GS
Gloss Selection
GR
Gloss Reordering
GT
ground truth
HMM
Hidden Markov Model
HPE
Hand Pose Enhancer
HOH
Hard of Hearing
HamNoSys
Hamburg Notation System
ISL
Irish Sign Language
LSTM
Long Short-Term Memory
MHA
Multi-Headed Attention
MoCap
Motion Capture
MTC
Monocular Total Capture
MSE
Mean Squared Error
MDN
Mixture Density Network
mDGS
Meine DGS Annotated
mDGST
meineDGST
mDGST-H
meineDGST-HARD
mDGST-E
meineDGST-EASY
MT
Machine Translation
NMT
Neural Machine Translation
NLP
Natural Language Processing
NAR
Non-AutoRegressive
NSVQ
Noise Substitution Vector Quantization
PHOENIX12
RWTH-PHOENIX-Weather-2012
PHOENIX14
RWTH-PHOENIX-Weather-2014
PHOENIX14T
RWTH-PHOENIX-Weather-2014T
POF
Part Orientation Field
POS
Part Of Speech
PT
Progressive Transformer
PAF
Part Affinity Field
P2TT
Pose-to-Text Transformer
P2T
Pose-to-Text
P2S
Pose-to-Sign
P2G2T
Pose-to-Gloss-to-Text
PCA
Principal Component Analysis
RELU
Rectified Linear Units
RNN
Recurrent Neural Network
ROUGE
Recall-Oriented Understudy for Gisting Evaluation
VQ-VAE
Vector Quantized Variational Autoencoders
VQ
Vector Quantisation
VAE
Variational Autoencoders
SGD
Stochastic Gradient Descent
SLA
Sign Language Assessment
SLR
Sign Language Recognition
SLT
Sign Language Translation
SLP
Sign Language Production
SMT
Statistical Machine Translation
SLO
Spoken Language Order
SO
Sign Language Order
S&R
Select and Reorder
SSE
Sign Supported English
S2T
Sign-to-Text
S2G2T
Sign-to-Gloss-to-Text
T2GT
Text-to-Gloss Transformer
T2PT
Text-to-Pose Transformer
T2P
Text-to-Pose
T2G
Text-to-Gloss
T2H
Text-to-HamNoSys
T2G2H
Text-to-Gloss-to-HamNoSys
T2G2P
Text-to-Gloss-to-Pose
T2S
Text-to-Sign
T2SSE
Text to Sign Supported English
WER
Word Error Rate
WMT2014
WMT2014 German-English
(eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version
11institutetext: CVSSP, University of Surrey
Guildford, United Kingdom
11email: {harry.walsh, b.saunders, r.bowden}@surrey.ac.uk
Sign Stitching: A Novel Approach to Sign Language Production
Harry Walsh    Ben Saunders    Richard Bowden
Abstract
Sign Language Production (SLP) is a challenging task, given the limited resources available and the inherent diversity within sign data. As a result, previous works have suffered from the problem of regression to the mean, leading to under-articulated and incomprehensible signing. In this paper, we propose using dictionary examples and a learnt codebook of facial expressions to create expressive sign language sequences. However, simply concatenating signs and adding the face creates robotic and unnatural sequences. To address this we present a 7-step approach to effectively stitch sequences together. First, by normalizing each sign into a canonical pose, cropping, and stitching we create a continuous sequence. Then, by applying filtering in the frequency domain and resampling each sign, we create cohesive natural sequences that mimic the prosody found in the original data. We leverage a SignGAN model to map the output to a photo-realistic signer and present a complete Text-to-Sign (T2S) Sign Language Production (SLP) pipeline. Our evaluation demonstrates the effectiveness of the approach, showcasing state-of-the-art performance across all datasets. Finally, a user evaluation shows our approach outperforms the baseline model and is capable of producing realistic sign language sequences.
Keywords:
Sign Language Production Continuous Sequence Synthesis Human Pose Generation
1 Introduction
Sign Language Production (SLP) is an essential step in facilitating two-way communication between the Deaf and Hearing communities. Sign language is inherently multi-channelled, with channels performed asynchronously and categorised into manual (hands and body) and non-manual (facial, rhythm, stress and intonation) features. For sign language to be truly understandable, both manual and non-manual features must be present.
Analogous to the tone and rhythm used in spoken language, signed language exhibits prosody. The natural rhythm, stress and intonation that signed languages use to convey additional information [50]. Other phenomena, such as the impact of sentiment on the manual and non-manual features have also been studied, showing that emotion can change the form of a sign to exhibit distinctive profiles [35].
Sign language corpora containing linguistic annotation are limited due to the cost and time required to create such annotations [23]. Previous works have attempted to directly regress a sequence of poses from the spoken language or representations such as gloss [36, 39, 14, 15, 52]. However, given that sign language is a low-resource language and the complexity is under-represented in small datasets, previous approaches have suffered from regression to the mean, resulting in under-articulated and incomprehensible signing. Additionally, previous works have implicitly modelled prosody, but due to the limited resources, it is often lost in production.
Figure 1: System overview of our approach to SLP. a) Spoken language to gloss, duration, facial token and cutoff Transformer. b) Stitcher that creates a pose sequence given a list of glosses, facial expressions, durations and a cutoff. c) The SignGAN module that produces a photo-realistic signer conditioned on a pose sequence.
In this paper, we propose a novel approach to SLP that effectively stitches together dictionary examples to create a meaningful, continuous sequence of sign language. By using isolated signs, we ensure the sequence remains expressive, overcoming previous shortcomings related to regression to the mean. However, each example lacks non-manual features, so we propose a Noise Substitution Vector Quantization (NSVQ) transformer architecture to learn a dictionary of facial expressions that can be added to each sign to create a realistic sequence. To the best of our knowledge, we are the first to explicitly model aspects of signed prosody in the context of SLP. By training a translation model to predict glosses, alongside a duration, facial expression and a cutoff value, we can modify the sequence to eliminate robotic and unnatural movements. Resampling each sign to the predicted duration allows us to alter the velocity associated with signing stress and rhythm [49]. Furthermore, by applying filtering in the frequency domain, we can adjust the trajectory of each sign to create softer signing, akin to how signers modify a sign to convey sentiment [3, 13]. Our approach demonstrates it is capable of modifying the stitched sequence to emulate aspects of prosody seen in the original continuous data. Evaluation of the produced sequence with back translation showcases state-of-the-art performance on all three datasets.
Furthermore, to conduct a realistic user evaluation we use SignGAN, a Generative Adversarial Network (GAN) capable of generating photo-realistic sign language videos from a sequence of poses [37]. Thus, we present a full Text-to-Sign (T2S) SLP pipeline that contains both manual and non-manual features. The user evaluation agrees that the approach outperforms the baseline method [38] and improves the realism of the signed sequence. An overview of the approach can be seen in Fig. 1.
Finally, when ground truth gloss timing is not available, such as on the RWTH-PHOENIX-Weather-2014T (PHOENIX14T) dataset, we leverage our stitching approach to perform sign segmentation. By creating an equivalent sequence and aligning it to the original data, we infer the duration of each sign.
The contributions of this paper can be summarised as:
  1. 1.
    A complete SLP pipeline, capable of generating expressive natural sign language videos from spoken language text.
  2. 2.
    A novel sign stitching approach to effectively join isolated examples and modify them to mimic sign prosody.
  3. 3.
    An approach to create a dictionary of non-manual facial features, that improves the realism of the produced sequence.
  4. 4.
    A novel approach to sign segmentation.
  5. 5.
    Comprehensive evaluation across several metrics and an in-depth user evaluation.
2 Related Work
2.0.1 Sign language Translation:
For the last 30 years Computational Sign Language Translation (SLT) has been an active area of research [44]. Initially focusing on isolated Sign Language Recognition (SLR) where a single instance of a sign was classified using a Convolutional Neural Network (CNN) [24]. Subsequent works extended to Continuous Sign Language Recognition (CSLR), which requires both the segmentation of a video into its constituent signs and their respective classification [22]. Later Camgoz et al. introduced the task of Sign-to-Text (S2T) [5] using neural networks, an extension of CSLR that requires the additional task of translation to spoken language. S2T performance was later improved using a Transformer [47]. By using gloss as a supervision on the encoder with a Connectionist Temporal Classification (CTC) loss the model achieved state-of-the-art performance [6]. Although there has been a lot of work since, the architecture has since become the standard when computing back-translation performance [36, 38, 39].
2.0.2 Sign Language Production:
SLP is the reverse task to SLT, which aims to translate spoken sentences into continuous sign language. Early approaches to SLP used an animated avatar driven with either motion capture data or parameterised glosses [2, 8, 53, 9, 46, 10]. These works all required expensive annotation systems, such as the Hamburg Notation System (HamNoSys) [33] or SigML [18] and have shown to be unpopular with the Deaf community due to the robotic motion and under articulated signing [34]. None of these approaches attempt to join the isolated signs effectively. Instead opting to play each sign in sequence, with unnatural transitions in between. Furthermore, due to the effect called the ‘uncanny valley’, users do not feel comfortable with the resulting production [28]. Methods that use Motion Capture (MoCap) data achieve better results, but have a limited vocabulary due to the cost of collecting and annotating data [11]. Here, we directly extract skeletons from sign language videos, overcoming the need for expensive MoCap. While Mediapipe has become a standard for this task [26], it sometimes provides incorrect keypoint predictions, negatively impacting downstream tasks. To enhance the quality of pose estimation and uplift the prediction to 3D, Ivashechkin et al. proposed a method that combines Forward Kinematics with a neural network to ensure valid predictions [16]. In this work, we use this additional optimization to extract an accurate sign dictionary.
Research suggests that over 15 million sentence pairs are necessary for Neural Machine Translation (NMT) to outperform statistical approaches [21], but sign Language datasets with parallel text-gloss pairs are restricted to 50,000 sequences [23]. Consequently, sign languages can be classified as a low-resource language, making SLP a challenging task.
Early deep learning SLP approaches used NMT and broke the task down into three steps, Text-to-Gloss (T2G), Gloss-to-Pose (G2P) and Pose-to-Sign (P2S) [43]. Saunders et al. introduced the Progressive Transformer (PT) [38], a transformer architecture that synthesises poses directly from text. Although better results were achieved using gloss as an intermediate representation, the approach suffered from regression to the mean, caused by the lack of training data and the diversity of lexical variants. To reduce the problem, adversarial training and Mixture Density Network (MDN) were applied [36, 39] and since then a range of approaches have been proposed [14, 15, 52]. However, visual inspection of the results shows that the approaches still suffer from regression to the mean, and as a result, they fail to effectively convey the translation. Here we propose a method to effectively join isolated signs, that does not suffer from regression to the mean. By filtering the movement in the frequency domain we ensure the produced sequence is stylistically cohesive. In addition, by modifying the length of each sign we recreate the prosody seen in the original data.
2.0.3 Vector Quantized Models:
The first Variational Autoencoders (VAE) model was introduced by Kingma et al. [20], but it struggled to capture fine-grain details. Later the first Vector Quantized Variational Autoencoders (VQ-VAE) architecture was introduced [30]. The model forced the latent space of the VAE to be discrete, and as a result, the model showed state-of-the-art image and audio generation performance. The model uses straight-though gradient estimation to propagate the gradient from the decoder to the encoder. This is necessary as the Vector Quantisation (VQ) operation uses an argmin operation to choose the closest matching codebook entry during training, and hence is non-differentiable. In total, the architecture required three losses to train. Kaiser et al. [17] reduced the required loss to two by adding an exponential moving average to update the codebook. Recently, Vali et al. [45] introduced the NSVQ VAE, this architecture further reduced the required loss functions to one. The technique approximated the quantization error and substituted it with normalised noise plus the original error. The model is trainable end-to-end, so we apply this approach to learn a dictionary of facial expressions.
2.0.4 Human Pose Synthesis:
Research indicates a strong preference among Deaf participants for sign language videos over skeleton representations [48]. Avatar approaches, due to their cartoon-like appearance, are also found to be unpopular, appearing more suitable for children rather than adults [51]. As a result, the field progressed to using photo-realistic approaches. Sanders et al. proposed the SignGAN architecture, a model capable of producing photo-realistic sign language videos, that are comprehensible by a native Deaf signer [37].
The model is designed as a GAN, trained with a combination of 5 losses. Once trained the generator network can synthesise pose and style conditioned images. We apply this model to conduct a user evaluation.
3 Methodology
SLP aims to facilitate the continuous translation from spoken to signed languages by converting a source spoken language sequence, $X=(x_{1},x_{2},...,x_{W})$ with W words into a video of photo-realistic sign, denoted as $V=(v_{1},v_{2},\ldots,v_{U})$ with U frames. To accomplish this we use two intermediate representations, following Fig. 1 from left to right. First, the spoken language is translated to a sequence of glosses, $Y=(y_{1},y_{2},\ldots,y_{G})$, face tokens, $\mathcal{F}^{t}=(f^{t}_{1},f^{t}_{2},\ldots,f^{t}_{G})$ and duration’s, $D=(d_{1},d_{2},\ldots,d_{G})$, all with length $G$. Additionally, for each sequence, we predict a low pass cutoff, $\mathcal{C}$ (Fig. 1.a). Each gloss and facial expression is stitched together using these parameters, to produce a continuous sequence of poses, denoted as $P=(p_{1},p_{2},...,p_{U})$ with U frames (Fig. 1.b). Finally, we use the pose sequence to condition the SignGAN module allowing us to produce a photo-realistic signer. Next, we provide a detailed explanation of each step in our pipeline, following the order illustrated in Fig. 1 from left to right. We then elaborate on the process of generating the cutoffs, gloss timestamps and the dictionary of facial expressions.
3.0.1 Translation Model
Given a spoken language sequence $X=(x_{1},x_{2},...,x_{W})$, our goal is to generate a corresponding sequence of glosses $Y=(y_{1},y_{2},...,y_{G})$. We design the transformer with four output layers, enabling the model to predict the corresponding duration (in frames) and facial expression for each gloss, plus a low-pass filter cutoff in Hz for each sequence. Thus the model learns the conditional probability $p(Y,D,\mathcal{F}^{t},C|X)$.
Figure 2: An overview of the Translation module.
The model is an encoder-decoder transformer with Multi-Headed Attention (MHA). The spoken language and gloss sequences are tokenized at the word level, and the embedding for a sequence is generated using a token embedding layer. Following the embedding layer, we add sine and cosine positional encoding.
The encoder learns to generate a contextualized representation for each token in the sequence. This representation is then fed into the decoder, which consists of multiple layers of self and cross MHA along with feedforward layers, and residual connections. The gloss, facial expression and duration predictions are obtained by passing the decoder output through their respective output layers. To obtain the cutoff prediction, we pool the decoder embedding across each time step and pass the output through a linear layer. The model is trained end-to-end with the following loss function;
$L_{total}=\lambda_{y}\sum_{i=1}^{Y}\hat{y}_{i}\log(y_{i})+\lambda_{f}\sum_{i=1}^{\mathcal{F}}\hat{F}_{i}\log(F_{i})+\lambda_{d}\frac{1}{n}\sum_{i=1}^{n}(d_{i}-\hat{d}_{i})^{2}+\lambda_{C}(C-\hat{C})^{2}$ (1)
Each component is scaled by a factor, $\lambda_{y}$, $\lambda_{d}$, $\lambda_{f}$ and $\lambda_{C}$ before being combined to give the total loss, $L_{total}$. The predictions from this model are passed to the stitching module to generate a pose sequence.
3.0.2 Stitching
For each dataset, we collect an isolated instance of each gloss in our target vocabulary. For each sign, we extract Mediapipe skeletons [26] and run an additional optimization to uplift the 2D skeletons to 3D [16]. The optimisation uses forward kinematics and a neural network to solve for joint angles, $J_{a}$. We choose to store our dictionary as joint angles, as this allows us to apply a canonical skeleton. This ensures the stitched sequence is consistent even if the original signers have different bone lengths. We define a dictionary of, $N_{s}$, signs as $DS=[S_{1},S_{2},...,S_{N_{s}}]$ where each sign in the dictionary consists of a sequence of angles, such that $S_{i}=(a_{1},a_{2},...,a_{U_{s}})$ and $a_{i}\in\mathbb{R}^{J_{a}}$, where $U_{s}$ is the duration in frames. In addition we define a learnt dictionary of, $N_{f}$, facial expressions as $DF=[F_{1},F_{2},...,F_{N_{f}}]$, where $F_{i}\in\mathbb{R}^{U_{f}\times J\times D}$. As illustrated in Fig. 3, the stitching pipeline is comprised of seven steps, we now detail each step.
Figure 3: An overview of the stitching module.
Step 1) Given a list of glosses, $Y$, we select the corresponding signs in our dictionary. If a gloss is absent from the dictionary, we initially lemmatize and format the gloss. If still, we are unable to find a match in the dictionary, we apply a word embedding model and compute the cosine similarity with all words in the dictionary. We then select the closest sign as the substitute. Such that;
$j_{sub}=\arg\max_{j}(\frac{\sum_{j=1}^{N_{s}}\varepsilon(y_{q})\cdot\varepsilon(DS^{y}_{j})}{\sqrt{\varepsilon(y_{q})^{2}}\cdot\sqrt{\sum_{j=1}^{N_{s}}\varepsilon(DS^{y}_{j})^{2}}})$ $S=DS[{j_{sub}}]$ (2)
Here $\varepsilon()$ represents the word embedding model, $y_{q}$ is the query gloss and $DS^{y}$ is the dictionary’s corresponding gloss tags. We find word embeddings capture the meaning of words, enabling substitutions such as replacing RUHRGEBIET (RUHR AREA) with LANDSCHAFT (LANDSCAPE). Simultaneously, in this step, we select the corresponding facial expressions, $F$, from the dictionary, $DF$, given the predicted face tokens, $\mathcal{F}^{t}$.
Step 2) The selected signs are converted from angles into a 3D canonical pose. We normalise the rotation of the signer, such that the midpoint of the hips is located at the origin and the shoulders are fixed on the y plane. This ensures the skeleton is consistent across all the signs. Consequently, we convert from a sequence of angles $S_{n}\in\mathbb{R}^{U_{\mathbf{s}}\times J_{a}}$ to a sequence of poses, $P=(p_{1},p_{2},...,p_{U_{s}})$ with the same number of frames, $U_{s}$. Each pose, $p_{u}$, is represented in $D$-dimensional space and consists of $J$ joints, denoted as $p_{u}\in\mathbb{R}^{J\times D}$.
Step 3) The dictionary signs often start and end from a rest pose. Therefore, to avoid unnatural transitions we cropped the beginning and end of each sign. For this, we track the keypoint $T$ corresponding to the wrist of the signer’s dominant hand and measure the distance travelled.
Thus, for each dictionary sign we create a sequence, $P^{\Delta}=(p^{\Delta}_{2},p^{\Delta}_{3},\ldots,p^{\Delta}_{U_{s}})$ $n\in{2,3,\ldots,U_{s}}$, representing the distance travelled for a dictionary sign. We remove the beginning frames once the sign has moved by a set threshold, $\alpha_{crop}$. The crop index is given by:
$\text{index}_{\text{start}}=\arg\max_{u}\left(\sum_{u=1}^{U_{s}}P^{\Delta}_{u}-\alpha_{crop}\cdot\sum_{u=1}^{\max(U_{s})}P^{\Delta}_{u}\right)$ (3)
To crop the end, we reverse the order of frames and repeat the process. However, for short sequences, this method may over-crop the sign, rendering it meaningless. We introduce an additional constraint $index_{end}-index_{start}<U_{min}$. If this condition holds, we reduce the threshold, $\alpha_{crop}$, by a factor of 0.1 and recalculate.
Step 4) As detailed above, we predict the duration of each gloss. Here, we utilize the duration to resample the length of each sign and facial expression, emulating the natural rhythm in the original data. This process involves upsampling or downsampling the sign using linear interpolation. However, due to co-articulation in the continuous data, some signs are predicted to be only a few frames long, which may not be sufficient for an isolated example. To address this, we impose a minimum sign length to ensure the produced sequence remains expressive. Once the facial expression and sign are resampled to the same length, we shift and rotate the face onto the body creating the complete skeleton.
Step 5) Having created a list of signs in a canonical pose, cropped, and resampled to match the original data, the next step involves joining these signs into a single sequence using the smart stitching module. The objective is to achieve a natural transition between the end of one sign and the start of the next. To accomplish this, we track the dominant hand of the signer and calculate the distance, $\Delta$, between the end of the first sign and the start of the second sign. Then we can determine the required number of frames, $U_{stitch}$, needed to create a smooth transition. Such that:
$U_{stitch}=\arg\min_{u}(V_{min}<\frac{\text{fps}*\Delta}{u}<V_{max})$ (4)
Where $V_{max}$ and $V_{min}$ are the start and end velocities of the two signs. This calculation ensures that the signer’s velocity is bounded by the end velocity of sign one and the initial velocity of sign two. In cases where multiple solutions exist, we select $U_{stitch}$ that minimizes the standard deviation between the start and end velocity. Following this, we employ linear interpolation to generate the missing frames.
Step 6) We concatenate the signs and the stitched frames to form a single sequence. We then sum all the predicted durations and resample the sequence so it matches the ground truth.
Step 7) Finally, we apply a low-pass Butterworth filter to each keypoint over time [4]. The predicted cutoff value determines which frequencies are removed, and corresponds to the -3dB attenuation point. This step aims to enhance the stylistic cohesiveness of the sequence by smoothing out any sharp, quick movements not present in the original sequence. The transfer function can be formulated as;
$H(z)=\frac{1}{{1+\left(\frac{z}{{e^{j\cdot\omega_{c}}}}\right)^{2N}}}$ (5)
Here we apply a 4th order filter thus, N is 4, and the angular cutoff frequency is given by $\omega_{c}=2\pi C$. z corresponds to the z-transform of the pose sequence. Applying the bilinear transform to Eq. 5 gives the discrete formula that we apply. This process generates natural pose sequences, that remain expressive and are stylistically cohesive. Next, we map these poses to a photo-realistic signer.
3.0.3 SignGAN
Skeleton outputs have been shown to reduce Deaf comprehension compared to a photo-realistic signer [48]. Therefore, to gain valuable feedback from the Deaf community we train a SignGan model [37].
Given a pose sequence, $P=(p_{1},p_{2},\ldots,p_{U})$, generated by our stitching approach the model aims to generate the corresponding video of sign language, $V=(v_{1},v_{2},\ldots,v_{U})$ with $U$ frames. Each video frame is conditioned on the appearance of an individual extracted from a style image. For the user evaluation, we use the appearance of a single individual obtained by passing a style image through the style encoder.
3.0.4 Facial Expression Generation
To be truly understandable and accepted by the Deaf community non-manual features must be present in the final output. Here we are using a discrete sequence-to-sequence model to generate the translation. Therefore, we must learn a discrete vocabulary of facial expressions, $DF$, that can be added to the isolated signs. We apply a NSVQ to learn a spatial-temporal dictionary of facial expressions.
For each gloss in the original data, we extract the corresponding face mesh, denoted as $F$. We then resample each sequence to a constant length, $U_{f}$ and scale it to be a constant size. Signers in the dataset are often looking off center, therefore we normalize the average direction of the face so that it is looking directly forward. Similar to Fig. 2 (Encoder) we add positional encoding and then embed the sequence using a single linear layer. After the embedding is passed through the transformer encoder to the codebook. The NSVQ codebook learns a set of $N_{f}$ embeddings. We denote the embedded face sequence and therefore each codebook entry as $F^{z}_{i}\in\mathbb{R}^{U_{f}\times H}$, where $H$ is the embedding dimension. Each input is mapped to one codebook entry, the difference between the selected codebook entry and the input is then simulated using a normally distributed noise source. A product of the simulated noise and the encoder output is then passed to the decoder. We use the counter decoding technique from the PT [38], to drive the decoder. The decoder learns to reconstruct the original face sequence and the input counter values. Thus, the model is trained with the following loss function;
$L_{Face}=\dfrac{1}{U_{f}}\sum_{u=1}^{U_{f}}((f_{u}-\hat{f_{u}})^{2}+\lambda_{CN}(c_{u}-\hat{c_{u}})^{2})$ (6)
Where $\lambda_{CN}$ is a scaling factor and $c$ is the counter value. Once fully trained we pass each codebook embedding, $F^{z}_{i}$, through the decoder to give the learnt dictionary of facial expressions in Euclidean space, $DF=[F_{1},F_{2},...,F_{N_{f}}]$.
3.0.5 Duration Generation
When ground truth timing information is not available we propose a novel sign segmentation approach based on the stitching method described above. Given the ground truth gloss labels, we generate the stitched sequence, $P_{stitch}$, but without step 4 (sign resampling).
Comparing the stitch sequence and the ground truth, we find that the motion can vary due to different lexical variants present compared to our dictionary. However, we find that the handshape is often still consistent. So, we take the keypoints that correspond to the signer’s hands and normalise the rotation so that the index finger metacarpal bone is fixed on the y-axis and the palm is fixed on the xy-plane, giving $P^{H}_{stitch},P^{H}$. Our next step is to align the two sequences so that we can infer the duration of the signs in the ground truth. For this we apply Dynamic Time Warping (DTW), such that;
$A_{i,j}=DTW(P^{H}_{stitch},P^{H})$ (7)
As we know the duration of the isolated signs in the stitched sequence, by analysing the alignment path, $A_{j}$, we can infer the duration of the signs in the original ground truth sequence.
3.0.6 Cutoff Generation
Experiments reveal that each sequence contains a distinct range of frequencies correlated to the signer’s style. Fast motions contain high frequencies, while soft, slow signing involves lower frequencies, typically within the range of 1 to 25 Hz.
To generate the ground truth cutoffs used in training, we once again apply our stitching approach. For each sequence in the ground truth data, $P$, we produce the equivalent stitched sequence, $P_{stitch}$. We then apply a low-pass filter to $P_{stitch}$ within the range of 1 to 25 Hz and measure the intersection and set difference of the frequencies, denoted as $(FT(P)\cap FT(P_{stitch}))$ and $(FT(P)\setminus FT(P_{stitch}))$, where the Fourier transform is represented as $FT()$. Subsequently, we fit a parametric spline curve to the intersection and set difference. To determine the cutoff we find the frequency that maximises the intersection while minimising the set difference. This provides the cutoff frequency for that sequence. We opt to use this method over just analyzing the frequency in the original sequence as we do not have an ideal filter. Thus, the butterworth filter has unintended effects on the frequency and phase blow the cutoff.
4 Experimental Setup
4.0.1 Implementation Details
In our experiments, we conducted a grid search for optimal hyper-parameters and identified the following settings as the most effective. Our encoder-decoder translation model is constructed with an embedding size of 512 and a feed-forward size of 1024. We find that the optimal number of layers and heads is dataset-dependent, with smaller datasets requiring fewer layers compared to Meine DGS Annotated (mDGS) where we use 3 layers and 4 heads. The models utilise dropout with a probability of 0.1 [41], ReLU activations between layers [1], and pre-layer normalisation for regularisation and training stability. Training employs a ’reduce on plateau’ scheduler with a patience of 5 and a decrease factor of 0.8. The layers are initialised using a Xavier initializer [12] with zero bias, and during training, Adam optimization is employed [19]. The initial learning rate is set to $10^{-4}$, and we train the model until convergence. During decoding, we utilise a greedy search algorithm. The loss scaling factors, $\lambda_{y}$, $\lambda_{d}$, $\lambda_{f}$ and $\lambda_{C}$ are set to 1.0, 0.1, 0.3 and 0.2, respectively. When stitching we enforce a minimum sign length, $U_{min}$ of 4 frames. The cropping threshold, $\alpha_{crop}$ is dataset-dependent, we find values between the range of 0.1 to 0.35 most effective. All sequences are subsampled to 12 frames per second (fps) for computational efficiency.
For each dataset, we create a dictionary of 500 facial expressions. We scale the counter loss by 0.01 and we set an embedding dimension of 512. The encoder and decoder are initialized with the same settings as our translation model.
The angular pose representation comprises 104 angles, while the Euclidean representation consists of 61 keypoints (21 for each hand and 19 for the body and face). The face mesh we add includes 128 keypoints, which are a subset of Mediapipe’s 478-face mesh [26].
For comparison, we train a progressive transformer on each dataset until convergence using the parameters from [38].
4.0.2 Datasets
The approach is tested on three datasets, the Public Corpus of German Sign Language, 3rd release, the mDGS dataset [23], PHOENIX14T [5] and the BSL Corpus T (BSLCPT) [40]. BSLCPT contains 211 participants from 8 regions in the UK, performing 4792 individual signs from a range of age groups. The participants perform narrative, interviews and participate in free conversation. Similarly, mDGS contains 330 participants engaging in free-form signing. Whereas, the PHOENIX14T dataset is extracted from German TV weather broadcasters and contains over 8,000 parallel sequences.
4.0.3 Dictionary
In the following experiment, we test two different dictionaries: 1) collected from isolated examples, and 2) a dictionary created from continuous data. Next, we provide further details about each:
Isolated: Here, the signs are sourced from individuals who perform each sign in isolation, typically starting from and returning to a resting position. When experimenting on the BSLCPT we use the Signbank dataset [7], it contains over 3,000 signs and includes all the lexical variants found in the BSLCPT dataset. However no such dictionary exists for the PHOENIX14T and mDGS dataset, therefore we collect a dictionary from a range of sources such as [23]. We find the mDGS dataset has a target vocabulary of 10,801. However, without the gloss variant, we find the core gloss vocabulary reduces to 4,434. We collect a total of 7,206 signs to experiment with. We use the method described in Sec. 3 (Stitching), step 1 to overcome issues with an incomplete vocabulary. To create word embeddings we apply Facebook’s implementation of Fasttext [27]. When experimenting on the BSLCPT we use the English implementation whereas on PHOENIX14T and mDGS we use the German implementation.
Continuous: Here we create a dictionary using the gloss timing annotations. The signs are taken from the test and dev data only, so that the back translation model has not seen the signs during training. As the examples come from the continuous sign, we omit the cropping step of the stitching pipeline. These dictionaries have an abundance of signs to choose from when stitching. We filter the dictionary and remove short signs as these are most likely co-articulated and therefore not suitable or out of context. We opt to randomly select the first sign in the sequence, and subsequent signs are chosen to ensure the most natural transition. Therefore, we select the sign from the dictionary in which the location of the wrist is closest to the last frame of the previous sign.
4.0.4 Evaluation Metrics
To evaluate our approach we employ a CSLR model (Sign Language Transformers [6]) to conduct back-translation, the same as [14, 36, 39, 52]. A model is trained for each dataset and the parameters are frozen so that results are consistent. BLEU [31], Rouge [25], and chrF [32] scores are computed between the predicted text and the ground truth. To evaluate the pose we employ Dynamic Time Warping Mean Joint Error (DTW-MJE), the metric compresses and stretches time to find the best alignment between the ground truth, $p$, and the predicted pose, $\hat{p}$.
5 Experiments
5.1 Quantitative Evaluation
5.1.1 Text-to-Gloss Translation Results
We start by evaluating the T2G translation performance described in Sec. 3.0.1. Tab. 1 shows the performance on all three datasets. We suggest that the difficulty of a dataset is proportional to the vocabulary and the total number of sequences used in training. We find the best performance on PHOENIX14T data which has the highest number of sequences per token, achieving 18.11 BLEU-4. In comparison to previous works, we find by having the model predict duration, face and cutoff we can achieve higher BLEU-1 scores, but at the cost of a lower BLEU-4 in comparison to [38]. On the more challenging mDGS dataset we find a considerably lower BLUE-4 score due to the larger domain of discourse. The BSLCPT has a smaller domain of discourse in comparison to mDGS but has the fewest sequences per token. Thus, understandably we only achieve a BLEU-4 of 1.67 on the test set. Overall we find the model to be performing as expected.
Table 1: The results of translating from Text-to-Gloss on the BSL Corpus T, RWTH-PHOENIX-Weather-2014T and Meine DGS Annotated dataset.
TEST SET DEV SET
Dataset: BLEU-1 BLEU-2 BLEU-3 BLEU-4 chrF ROUGE BLEU-1 BLEU-2 BLEU-3 BLEU-4 chrF ROUGE
BSLCPT 26.02 11.19 4.15 1.67 23.54 25.06 26.88 11.40 4.72 1.28 23.59 26.96
mDGS 30.13 13.04 5.45 2.36 29.24 31.43 29.72 12.46 4.89 1.87 28.90 30.90
PHOENIX14T 55.48 36.54 25.18 18.11 49.30 53.83 56.55 37.32 25.85 18.74 48.91 54.81
PHOENIX14T [38] 55.18 37.10 26.24 19.10 - 54.55 55.65 38.21 27.36 20.23 - 55.41
PHOENIX14T [42] 50.67 32.25 21.54 15.26 - 48.10 50.15 32.47 22.30 16.34 - 48.42
5.1.2 Text-to-Pose Translation Results
Note in the following experiments the back-translation model’s performance (shown as GT top row of Tab. 2, 3 and 4) should be considered the upper limit of performance. In this section, we evaluate the Text-to-Pose (T2P) performance using back translation. To allow for a comparison we train two versions of the PT with the setting presented in [38]. PT is the standard architecture, while PT + GN is trained with Gaussian Noise added to the input. In line with the original paper, we find Gaussian Noise improves the performance, however, our approach still outperforms both models except on DTW-MJE. As discussed previously, other works suffer from regression to the mean caused by the models attempting to minimise their loss function and thus, are incentivised to predict a mean pose. This metric fails to evaluate the content of the sequence, but the higher score does indicate our model is expressive as it is producing sequences further from the mean. For back-translation, we outperform the PT on all metrics. Showing significant improvements in BLEU-1 score of 98% and 269% on the mDGS and BSLCPT dev set (comparing PT + GN and Stitcher (continuous), Tab. 2 and 3).
Deep learning models exhibit a bias toward the data they were trained on and often show poor out-of-domain performance. Unsurprisingly, the performance improves when using the continuous dictionary. We find only a small increase in BLEU-1 of 0.13 on the BSLCPT dev set (Tab. 2), most likely due to the isolated dictionary containing the lexical variants found in the original data. Whereas we see a larger increase on the mDGS dataset (Tab. 3).
Table 2: The results of translating from Text-to-Pose on the BSL Corpus T dataset.
BSLCPT TEST SET DEV SET
Approach: DTW-MJE BLEU-1 BLEU-2 BLEU-3 BLEU-4 chrF ROUGE DTW-MJE BLEU-1 BLEU-2 BLEU-3 BLEU-4 chrF ROUGE
GT 0.000 17.3 3.96 1.37 0.54 13.00 21.76 0.000 17.32 3.71 1.08 0.39 13.04 21.89
PT [38] 0.288 4.40 0.65 0.18 0.00 5.80 8.22 0.292 4.00 0.61 0.10 0.00 5.69 8.02
PT + GN [38] 0.267 4.96 0.55 0.13 0.00 6.38 8.82 0.258 4.47 0.63 0.09 0.00 6.14 8.89
Stitcher (Isolated) 0.588 16.37 2.86 0.75 0.28 14.07 20.84 0.592 16.39 2.82 0.58 0.00 13.9 19.55
Stitcher (continuous) 0.575 16.99 3.65 1.03 0.41 14.32 20.65 0.573 16.52 3.19 0.73 0.00 14.34 20.53
Table 3: The results of translating from Text-to-Pose on the Meine DGS Annotated (mDGS) dataset.
mDGS TEST SET DEV SET
Approach: DTW-MJE BLEU-1 BLEU-2 BLEU-3 BLEU-4 chrF ROUGE DTW-MJE BLEU-1 BLEU-2 BLEU-3 BLEU-4 chrF ROUGE
GT 0.000 20.87 5.60 1.89 0.80 17.56 23.78 0.000 20.75 5.43 1.81 0.76 17.63 23.41
PT [38] 0.229 6.11 0.94 0.21 0.05 8.07 8.36 0.228 6.22 0.98 0.17 0.00 8.23 8.44
PT + GN [38] 0.2245 7.18 1.48 0.40 0.01 8.46 8.38 0.2241 9.22 1.63 0.38 0.01 8.94 8.57
Stitcher (Isolated) 0.581 16.63 3.75 0.94 0.22 13.39 21.69 0.592 16.9 3.67 0.95 0.32 13.9 21.34
Stitcher (Continuous) 0.637 18.64 4.17 1.07 0.39 16.86 21.80 0.637 18.27 4.07 1.19 0.43 16.75 21.25
Previous work has primarily focused on G2P translation, therefore to facilitate a meaningful comparison we present two versions of the model. First, a G2P version, where we use the ground truth data and just apply the stitching module, and, secondly our T2P approach (translation then stitching). Results for comparison are provided by [52]. We find our approach outperforms previous work on the BLEU-1 to 2 scores increasing the score by 56% and 19%, respectively (comparing Tab. 4, row 7 and 11). We also find significant improvement in ROUGE and chrF metrics.
Table 4: The results of translating from Gloss-to-Pose (G2P) and Text-to-Pose (T2P) on the RWTH-PHOENIX-Weather-2014T dataset.
PHOENIX14T
Approach: DTW-MJE BLEU-1 BLEU-2 BLEU-3 BLEU-4 chrF ROUGE
GT 0.000 32.41 20.19 14.41 11.32 33.84 32.96
PT [38] 0.197 6.27 3.33 2.14 1.59 14.52 9.50
PT + GN [38] 0.191 11.45 7.08 5.08 4.04 19.09 14.52
NAT-AT [14] 0.177 14.26 9.93 7.11 5.53 21.87 18.72
NAT-EA [14] 0.146 15.12 10.45 7.99 6.66 22.98 19.43
PoseVQ-MP [52] 0.146 15.43 10.69 8.26 6.98 - -
PoseVQ-DDM [52] 0.116 16.11 11.37 9.22 7.50 - -
Stitching G2P (Isolated) 0.593 21.47 8.79 4.25 2.49 23.74 20.32
Stitching G2P (Continuous) 0.587 23.58 12.31 8.05 5.95 28.85 24.84
Stitching T2P (Isolated) 0.594 22.78 9.68 5.17 3.12 24.27 21.30
Stitching T2P (Continuous) 0.572 25.14 13.54 8.98 6.67 29.5 26.49
Using a continuous dictionary we can outperform all models except for the VQ based approaches on BLEU-3 to 4. As the VQ model is learning sub-units of a gloss sequence we suggest this gives it an advantage on higher n-grams, as each token can represent multiple signs.
5.1.3 Segmentation Results
To evaluate the stitching segmentation approach we calculate the duration for each gloss in the mDGS dataset test set. We achieve a sign level frame F1-score of 0.6373 a similar score compared to [29] that achieves a top score of 0.63. Demonstrating the effectiveness of stitching for sign segmentation. It is worth noting our approach requires gloss information, but is computationally inexpensive compared to the LSTM used in [29].
5.2 Qualitative Evaluation
5.2.1 Visual Outputs
To demonstrate the approach’s effectiveness, we present skeleton and video outputs for two sign languages (BSL and DGS) [^1] https://github.com/walsharry/Sign_Stitching_Demos. To enable a fair evaluation, we also share failure cases where incorrect predictions from the initial translation produce repetitive sequences. Furthermore, in the supplementary material, we share visualizations of the produced sequence.
5.2.2 Survey Results
The survey presented a comparison to PT, followed by an ablation of different components of the stitching approach. 17% of people surveyed were native Deaf signers, while 34% were L2 signers or language learners. 87.5% preferred our approach compared to the PT, while the rest selected no preference. 100% of people agreed that applying the filter improved the realism compared to no filtering, while resampling was found to be less important, with 37.5% selecting no preference between the resampled and normal sequence. Interestingly, 41.7% preferred using an isolated dictionary, while 20.8% preferred the continuous dictionary and the rest indicated no preference.
6 Conclusion
In this paper, we presented a novel approach to SLP. Previous works have suffered from the problem of regression to the mean and have mainly focused on manual features. Here we have overcome the problem by using a dictionary of expressive examples. The stitching effectively joins the signs together creating a natural continuous sequence and by clustering facial expressions into a vocabulary we can create a sequence that contains both manual and non-manual features. We eliminated unnatural transitions and enhanced the stylistic cohesiveness through the approach. As a result, we present state-of-the-art performance. Finally, the user evaluation agrees with the quantitative results, indicating our approach can produce realistic expressive Sign language.

### v2

@@
 WMT2014 German-English
-(eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version
-11institutetext: CVSSP, University of Surrey
-Guildford, United Kingdom
-11email: {harry.walsh, b.saunders, r.bowden}@surrey.ac.uk
+[…]
+[…]
+University of Surrey
+Guildford, UK Sign Stitching
 Sign Stitching: A Novel Approach to Sign Language Production
-Harry Walsh    Ben Saunders    Richard Bowden
 Abstract
@@
 As a result, previous works have suffered from the problem of regression to the mean, leading to under-articulated and incomprehensible signing.
-In this paper, we propose using dictionary examples and a learnt codebook of facial expressions to create expressive sign language sequences.
-However, simply concatenating signs and adding the face creates robotic and unnatural sequences.
-To address this we present a 7-step approach to effectively stitch sequences together.
-First, by normalizing each sign into a canonical pose, cropping, and stitching we create a continuous sequence.
-Then, by applying filtering in the frequency domain and resampling each sign, we create cohesive natural sequences that mimic the prosody found in the original data.
-We leverage a SignGAN model to map the output to a photo-realistic signer and present a complete Text-to-Sign (T2S) Sign Language Production (SLP) pipeline.
-Our evaluation demonstrates the effectiveness of the approach, showcasing state-of-the-art performance across all datasets.
-Finally, a user evaluation shows our approach outperforms the baseline model and is capable of producing realistic sign language sequences.
-Keywords:
-Sign Language Production Continuous Sequence Synthesis Human Pose Generation
+In this paper, we propose using dictionary examples to create expressive sign language sequences.
+However, simply concatenating the signs would create robotic and unnatural sequences.
+Therefore, we present a 7-step approach to effectively stitch the signs together.
+First, by normalising each sign into a canonical pose, cropping and stitching we create a continuous sequence.
+Then by applying filtering in the frequency domain and resampling each sign we create cohesive natural sequences, that mimic the prosody found in the original data.
+We leverage the SignGAN model to map the output to a photo-realistic signer and present a complete Text-to-Sign (T2S) SLP pipeline.
+Our evaluation demonstrates the effectiveness of this approach, showcasing state-of-the-art performance across all datasets.
 1 Introduction
@@
 For sign language to be truly understandable, both manual and non-manual features must be present.
-Analogous to the tone and rhythm used in spoken language, signed language exhibits prosody.
-The natural rhythm, stress and intonation that signed languages use to convey additional information [50].
-Other phenomena, such as the impact of sentiment on the manual and non-manual features have also been studied, showing that emotion can change the form of a sign to exhibit distinctive profiles [35].
-Sign language corpora containing linguistic annotation are limited due to the cost and time required to create such annotations [23].
-Previous works have attempted to directly regress a sequence of poses from the spoken language or representations such as gloss [36, 39, 14, 15, 52].
+Analogous to the tone and rhythm used in spoken language, signed language exhibits prosody. The natural rhythm, stress and intonation that signed languages use to convey information [Wilbur(2009)].
+Figure 1: Overview of our approach. a) Spoken language to gloss, duration, facial expression and cutoff translation. b) Pose sequence generation. c) Photorealistic signer production.
+Sign language corpora containing linguistic annotation are limited due to the cost and time required to create such annotations [Konrad et al.(2020)Konrad, Hanke, Langer, Blanck, Bleicken, Hofmann, Jeziorski, König, König, Nishio, Regen, Salden, Wagner, Worseck, Böse, Jahn, and Schulder].
+Previous works have attempted to directly regress a sequence of poses from the spoken language or representations such as gloss [Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden, Huang et al.(2021)Huang, Pan, Zhao, and Tian, Hwang et al.(2023)Hwang, Lee, and Park, Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu, Xie et al.(2024)Xie, Zhang, Taiying, Tang, Du, and Li, Stoll et al.(2020)Stoll, Camgoz, Hadfield, and Bowden, Tang et al.(2022)Tang, Hong, Guo, and Wang].
 However, given that sign language is a low-resource language and the complexity is under-represented in small datasets, previous approaches have suffered from regression to the mean, resulting in under-articulated and incomprehensible signing.
 Additionally, previous works have implicitly modelled prosody, but due to the limited resources, it is often lost in production.
-Figure 1: System overview of our approach to SLP.
-a) Spoken language to gloss, duration, facial token and cutoff Transformer.
-b) Stitcher that creates a pose sequence given a list of glosses, facial expressions, durations and a cutoff.
-c) The SignGAN module that produces a photo-realistic signer conditioned on a pose sequence.
 In this paper, we propose a novel approach to SLP that effectively stitches together dictionary examples to create a meaningful, continuous sequence of sign language.
@@
 By training a translation model to predict glosses, alongside a duration, facial expression and a cutoff value, we can modify the sequence to eliminate robotic and unnatural movements.
-Resampling each sign to the predicted duration allows us to alter the velocity associated with signing stress and rhythm [49].
-Furthermore, by applying filtering in the frequency domain, we can adjust the trajectory of each sign to create softer signing, akin to how signers modify a sign to convey sentiment [3, 13].
+Resampling each sign to the predicted duration allows us to alter the velocity associated with signing stress and rhythm [Wilbur(1999)].
+Furthermore, by applying filtering in the frequency domain, we can adjust the trajectory of each sign to create softer signing, akin to how signers modify a sign to convey sentiment [Brentari(1998), Heloir and Gibet(2007), Reilly et al.(1992)Reilly, McIntire, and Seago].
 Our approach demonstrates it is capable of modifying the stitched sequence to emulate aspects of prosody seen in the original continuous data.
 Evaluation of the produced sequence with back translation showcases state-of-the-art performance on all three datasets.
-Furthermore, to conduct a realistic user evaluation we use SignGAN, a Generative Adversarial Network (GAN) capable of generating photo-realistic sign language videos from a sequence of poses [37].
+Furthermore, to conduct a realistic user evaluation we use SignGAN, a Generative Adversarial Network (GAN) capable of generating photo-realistic sign language videos from a sequence of poses [Saunders et al.(2020b)Saunders, Camgoz, and Bowden].
 Thus, we present a full Text-to-Sign (T2S) SLP pipeline that contains both manual and non-manual features.
-The user evaluation agrees that the approach outperforms the baseline method [38] and improves the realism of the signed sequence.
+The user evaluation agrees that the approach outperforms the baseline method [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] and improves the realism of the signed sequence.
 An overview of the approach can be seen in Fig.
 1.
-Finally, when ground truth gloss timing is not available, such as on the RWTH-PHOENIX-Weather-2014T (PHOENIX14T) dataset, we leverage our stitching approach to perform sign segmentation.
-By creating an equivalent sequence and aligning it to the original data, we infer the duration of each sign.
-The contributions of this paper can be summarised as:
-  1. 1.
-    A complete SLP pipeline, capable of generating expressive natural sign language videos from spoken language text.
-  2. 2.
-    A novel sign stitching approach to effectively join isolated examples and modify them to mimic sign prosody.
-  3. 3.
-    An approach to create a dictionary of non-manual facial features, that improves the realism of the produced sequence.
-  4. 4.
-    A novel approach to sign segmentation.
-  5. 5.
-    Comprehensive evaluation across several metrics and an in-depth user evaluation.
 2 Related Work
-2.0.1 Sign language Translation:
-For the last 30 years Computational Sign Language Translation (SLT) has been an active area of research [44].
-Initially focusing on isolated Sign Language Recognition (SLR) where a single instance of a sign was classified using a Convolutional Neural Network (CNN) [24].
-Subsequent works extended to Continuous Sign Language Recognition (CSLR), which requires both the segmentation of a video into its constituent signs and their respective classification [22].
+Sign language Translation: For the last 30 years Computational Sign Language Translation (SLT) has been an active area of research [Tamura and Kawasaki(1988)].
+Initially focusing on isolated Sign Language Recognition (SLR) where a single instance of a sign was classified using a Convolutional Neural Network (CNN) [Lecun et al.(1998)Lecun, Bottou, Bengio, and Haffner].
+Subsequent works extended to Continuous Sign Language Recognition (CSLR), which requires both the segmentation of a video into its constituent signs and their respective classification [Koller et al.(2015)Koller, Forster, and Ney].
 Later Camgoz et al.
-introduced the task of Sign-to-Text (S2T) [5] using neural networks, an extension of CSLR that requires the additional task of translation to spoken language.
-S2T performance was later improved using a Transformer [47].
-By using gloss as a supervision on the encoder with a Connectionist Temporal Classification (CTC) loss the model achieved state-of-the-art performance [6].
-Although there has been a lot of work since, the architecture has since become the standard when computing back-translation performance [36, 38, 39].
-2.0.2 Sign Language Production:
-SLP is the reverse task to SLT, which aims to translate spoken sentences into continuous sign language.
-Early approaches to SLP used an animated avatar driven with either motion capture data or parameterised glosses [2, 8, 53, 9, 46, 10].
-These works all required expensive annotation systems, such as the Hamburg Notation System (HamNoSys) [33] or SigML [18] and have shown to be unpopular with the Deaf community due to the robotic motion and under articulated signing [34].
+introduced the task of Sign-to-Text (S2T) [Camgoz et al.(2018)Camgoz, Hadfield, Koller, Ney, and Bowden] using neural networks, an extension of CSLR that requires the additional task of translation to spoken language.
+S2T performance was later improved using a Transformer [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin].
+Although there has been a lot of work since, the architecture has since become the standard when computing back-translation performance [Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2020c)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden].
+Sign Language Production: SLP is the reverse task to SLT, which aims to translate spoken sentences into continuous sign language.
+Early approaches to SLP used an animated avatar driven with either motion capture data or parameterised glosses [Bangham et al.(2000)Bangham, Cox, Elliott, Glauert, Marshall, Rankov, and Wells, Cox et al.(2002)Cox, Lincoln, Tryggvason, Nakisa, Wells, Tutt, and Abbott, Zwitserlood et al.(2004)Zwitserlood, Verlinden, Ros, Van Der Schoot, and Netherlands, Efthimiou et al.(2012)Efthimiou, Fotinea, Hanke, Glauert, Bowden, Braffort, Collet, Maragos, and Lefebvre-Albaret, Van Wyk(2008), ElGhoul and Jemni(2011)].
+These works all required expensive annotation systems, such as the Hamburg Notation System (HamNoSys) [Prillwitz and für Deutsche Gebärdensprache und Kommunikation Gehörloser(1989)] or SigML [Kennaway(2015)] and have shown to be unpopular with the Deaf community due to the robotic motion and under articulated signing [Rastgoo et al.(2021)Rastgoo, Kiani, Escalera, and Sabokrou].
 None of these approaches attempt to join the isolated signs effectively.
-Instead opting to play each sign in sequence, with unnatural transitions in between.
-Furthermore, due to the effect called the ‘uncanny valley’, users do not feel comfortable with the resulting production [28].
-Methods that use Motion Capture (MoCap) data achieve better results, but have a limited vocabulary due to the cost of collecting and annotating data [11].
-Here, we directly extract skeletons from sign language videos, overcoming the need for expensive MoCap.
-While Mediapipe has become a standard for this task [26], it sometimes provides incorrect keypoint predictions, negatively impacting downstream tasks.
-To enhance the quality of pose estimation and uplift the prediction to 3D, Ivashechkin et al.
-proposed a method that combines Forward Kinematics with a neural network to ensure valid predictions [16].
-In this work, we use this additional optimization to extract an accurate sign dictionary.
-Research suggests that over 15 million sentence pairs are necessary for Neural Machine Translation (NMT) to outperform statistical approaches [21], but sign Language datasets with parallel text-gloss pairs are restricted to 50,000 sequences [23].
-Consequently, sign languages can be classified as a low-resource language, making SLP a challenging task.
-Early deep learning SLP approaches used NMT and broke the task down into three steps, Text-to-Gloss (T2G), Gloss-to-Pose (G2P) and Pose-to-Sign (P2S) [43].
+Instead opting to play each sign in the sequence, with unnatural transitions in between.
+Early deep learning SLP approaches used Neural Machine Translation (NMT) and broke the task down into three steps, Text-to-Gloss (T2G), Gloss-to-Pose (G2P) and Pose-to-Sign (P2S) [Stoll et al.(2020)Stoll, Camgoz, Hadfield, and Bowden].
 Saunders et al.
-introduced the Progressive Transformer (PT) [38], a transformer architecture that synthesises poses directly from text.
+introduced the Progressive Transformer (PT) [Saunders et al.(2020c)Saunders, Camgoz, and Bowden], a transformer architecture that synthesises poses directly from text.
 Although better results were achieved using gloss as an intermediate representation, the approach suffered from regression to the mean, caused by the lack of training data and the diversity of lexical variants.
-To reduce the problem, adversarial training and Mixture Density Network (MDN) were applied [36, 39] and since then a range of approaches have been proposed [14, 15, 52].
+To reduce the problem, adversarial training and Mixture Density Network (MDN) were applied [Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden] and since then a range of approaches have been proposed [Huang et al.(2021)Huang, Pan, Zhao, and Tian, Hwang et al.(2023)Hwang, Lee, and Park, Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu, Xie et al.(2024)Xie, Zhang, Taiying, Tang, Du, and Li, Tang et al.(2022)Tang, Hong, Guo, and Wang, Saunders et al.(2021a)Saunders, Camgoz, and Bowden].
 However, visual inspection of the results shows that the approaches still suffer from regression to the mean, and as a result, they fail to effectively convey the translation.
-Here we propose a method to effectively join isolated signs, that does not suffer from regression to the mean.
-By filtering the movement in the frequency domain we ensure the produced sequence is stylistically cohesive.
-In addition, by modifying the length of each sign we recreate the prosody seen in the original data.
-2.0.3 Vector Quantized Models:
-The first Variational Autoencoders (VAE) model was introduced by Kingma et al.
-[20], but it struggled to capture fine-grain details.
-Later the first Vector Quantized Variational Autoencoders (VQ-VAE) architecture was introduced [30].
-The model forced the latent space of the VAE to be discrete, and as a result, the model showed state-of-the-art image and audio generation performance.
-The model uses straight-though gradient estimation to propagate the gradient from the decoder to the encoder.
-This is necessary as the Vector Quantisation (VQ) operation uses an argmin operation to choose the closest matching codebook entry during training, and hence is non-differentiable.
-In total, the architecture required three losses to train.
-Kaiser et al.
-[17] reduced the required loss to two by adding an exponential moving average to update the codebook.
-Recently, Vali et al.
-[45] introduced the NSVQ VAE, this architecture further reduced the required loss functions to one.
-The technique approximated the quantization error and substituted it with normalised noise plus the original error.
-The model is trainable end-to-end, so we apply this approach to learn a dictionary of facial expressions.
-2.0.4 Human Pose Synthesis:
-Research indicates a strong preference among Deaf participants for sign language videos over skeleton representations [48].
-Avatar approaches, due to their cartoon-like appearance, are also found to be unpopular, appearing more suitable for children rather than adults [51].
-As a result, the field progressed to using photo-realistic approaches.
-Sanders et al.
-proposed the SignGAN architecture, a model capable of producing photo-realistic sign language videos, that are comprehensible by a native Deaf signer [37].
-The model is designed as a GAN, trained with a combination of 5 losses.
-Once trained the generator network can synthesise pose and style conditioned images.
-We apply this model to conduct a user evaluation.
+Here we propose a method to effectively join isolated signs, which means the produced sequences are guaranteed to be expressive and do not suffer from regression to the mean.
+Furthermore, preliminary experiments reveal that each sign language sequence contains a distinct range of frequencies correlated to the signer’s style.
+Fast motions contain high frequencies, while soft, slow signing involves lower frequencies, typically within the range of 1 to 25 Hz.
+To emulate this characteristic we filter the produced sequences in the frequency domain using a low-pass Butterworth filter [Butterworth et al.(1930)].
+This ensures the movements are stylistically cohesive.
+In addition, by adjusting the duration of each sign, we recreate the prosody observed in the original data.
 3 Methodology
 SLP aims to facilitate the continuous translation from spoken to signed languages by converting a source spoken language sequence, $X=(x_{1},x_{2},...,x_{W})$ with W words into a video of photo-realistic sign, denoted as $V=(v_{1},v_{2},\ldots,v_{U})$ with U frames.
-To accomplish this we use two intermediate representations, following Fig.
+To accomplish this the approach uses two intermediate representations, following Fig.
 1 from left to right.
 First, the spoken language is translated to a sequence of glosses, $Y=(y_{1},y_{2},\ldots,y_{G})$, face tokens, $\mathcal{F}^{t}=(f^{t}_{1},f^{t}_{2},\ldots,f^{t}_{G})$ and duration’s, $D=(d_{1},d_{2},\ldots,d_{G})$, all with length $G$.
-Additionally, for each sequence, we predict a low pass cutoff, $\mathcal{C}$ (Fig.
+Additionally, for each sequence, we predict a low pass cutoff, $\mathcal{C}$, which we use to filter the movements (Fig.
 1.a).
@@
 1 from left to right.
-We then elaborate on the process of generating the cutoffs, gloss timestamps and the dictionary of facial expressions.
-3.0.1 Translation Model
+We then elaborate on the process of generating the cutoff frequencies and the dictionary of facial expressions.
+3.1 Translation Model
 Given a spoken language sequence $X=(x_{1},x_{2},...,x_{W})$, our goal is to generate a corresponding sequence of glosses $Y=(y_{1},y_{2},...,y_{G})$.
 We design the transformer with four output layers, enabling the model to predict the corresponding duration (in frames) and facial expression for each gloss, plus a low-pass filter cutoff in Hz for each sequence.
-Thus the model learns the conditional probability $p(Y,D,\mathcal{F}^{t},C|X)$.
+Thus the model learns the conditional probability $p(Y,D,\mathcal{F}^{t},\mathcal{C}|X)$.
 Figure 2: An overview of the Translation module.
@@
 The predictions from this model are passed to the stitching module to generate a pose sequence.
-3.0.2 Stitching
+3.2 Stitching
 For each dataset, we collect an isolated instance of each gloss in our target vocabulary.
-For each sign, we extract Mediapipe skeletons [26] and run an additional optimization to uplift the 2D skeletons to 3D [16].
+For each sign, we extract Mediapipe skeletons [Lugaresi et al.(2019)Lugaresi, Tang, Nash, McClanahan, Uboweja, Hays, Zhang, Chang, Yong, Lee, et al.] and run an additional optimization to uplift the 2D skeletons to 3D [Ivashechkin et al.(2023)Ivashechkin, Mendez, and Bowden].
 The optimisation uses forward kinematics and a neural network to solve for joint angles, $J_{a}$.
@@
 In addition we define a learnt dictionary of, $N_{f}$, facial expressions as $DF=[F_{1},F_{2},...,F_{N_{f}}]$, where $F_{i}\in\mathbb{R}^{U_{f}\times J\times D}$.
-As illustrated in Fig.
-3, the stitching pipeline is comprised of seven steps, we now detail each step.
+Fig.
+3 illustrates our seven-step stitching pipeline, we now detail each step.
 Figure 3: An overview of the stitching module.
@@
 Thus, for each dictionary sign we create a sequence, $P^{\Delta}=(p^{\Delta}_{2},p^{\Delta}_{3},\ldots,p^{\Delta}_{U_{s}})$ $n\in{2,3,\ldots,U_{s}}$, representing the distance travelled for a dictionary sign.
-We remove the beginning frames once the sign has moved by a set threshold, $\alpha_{crop}$.
+We remove the beginning frames once the sign has moved by a threshold, $\alpha_{crop}$.
 The crop index is given by:
@@
 To crop the end, we reverse the order of frames and repeat the process.
-However, for short sequences, this method may over-crop the sign, rendering it meaningless.
-We introduce an additional constraint $index_{end}-index_{start}<U_{min}$.
-If this condition holds, we reduce the threshold, $\alpha_{crop}$, by a factor of 0.1 and recalculate.
 Step 4) As detailed above, we predict the duration of each gloss.
@@
 This process involves upsampling or downsampling the sign using linear interpolation.
-However, due to co-articulation in the continuous data, some signs are predicted to be only a few frames long, which may not be sufficient for an isolated example.
-To address this, we impose a minimum sign length to ensure the produced sequence remains expressive.
 Once the facial expression and sign are resampled to the same length, we shift and rotate the face onto the body creating the complete skeleton.
@@
 Following this, we employ linear interpolation to generate the missing frames.
-Step 6) We concatenate the signs and the stitched frames to form a single sequence. We then sum all the predicted durations and resample the sequence so it matches the ground truth.
-Step 7) Finally, we apply a low-pass Butterworth filter to each keypoint over time [4].
+Step 6) We concatenate the signs and the stitched frames to form a single sequence. We then sum all the predicted durations and resample the sequence to match the ground truth.
+Step 7) Finally, we apply a low-pass Butterworth filter to each keypoint over time [Butterworth et al.(1930)].
 The predicted cutoff value determines which frequencies are removed, and corresponds to the -3dB attenuation point.
@@
 The transfer function can be formulated as;
-$H(z)=\frac{1}{{1+\left(\frac{z}{{e^{j\cdot\omega_{c}}}}\right)^{2N}}}$ (5)
+$H(z)=\left((1+\left(\frac{z}{{e^{j\cdot\omega_{c}}}}\right)^{2N}\right)^{-1}$ (5)
 Here we apply a 4th order filter thus, N is 4, and the angular cutoff frequency is given by $\omega_{c}=2\pi C$.
@@
 Next, we map these poses to a photo-realistic signer.
-3.0.3 SignGAN
-Skeleton outputs have been shown to reduce Deaf comprehension compared to a photo-realistic signer [48]. Therefore, to gain valuable feedback from the Deaf community we train a SignGan model [37].
+3.3 SignGAN
+Skeleton outputs have been shown to reduce Deaf comprehension compared to a photo-realistic signer [Ventura et al.(2020)Ventura, Duarte, and Giró-i Nieto].
+Therefore, to gain valuable feedback from the Deaf community we train a SignGan model [Saunders et al.(2020b)Saunders, Camgoz, and Bowden].
 Given a pose sequence, $P=(p_{1},p_{2},\ldots,p_{U})$, generated by our stitching approach the model aims to generate the corresponding video of sign language, $V=(v_{1},v_{2},\ldots,v_{U})$ with $U$ frames.
-Each video frame is conditioned on the appearance of an individual extracted from a style image.
-For the user evaluation, we use the appearance of a single individual obtained by passing a style image through the style encoder.
-3.0.4 Facial Expression Generation
+3.4 Facial Expression Generation
 To be truly understandable and accepted by the Deaf community non-manual features must be present in the final output.
@@
 Therefore, we must learn a discrete vocabulary of facial expressions, $DF$, that can be added to the isolated signs.
-We apply a NSVQ to learn a spatial-temporal dictionary of facial expressions.
-For each gloss in the original data, we extract the corresponding face mesh, denoted as $F$.
+We design a transformer base NSVQ architecture to learn a spatial-temporal dictionary of facial expressions.
+Using the ground truth gloss timing information, we extract the corresponding face mesh sequence, denoted as $F$.
 We then resample each sequence to a constant length, $U_{f}$ and scale it to be a constant size.
@@
 After the embedding is passed through the transformer encoder to the codebook.
-The NSVQ codebook learns a set of $N_{f}$ embeddings.
+The NSVQ codebook learns a set of $N_{f}$ embeddings [Vali and Bäckström(2022)].
 We denote the embedded face sequence and therefore each codebook entry as $F^{z}_{i}\in\mathbb{R}^{U_{f}\times H}$, where $H$ is the embedding dimension.
@@
 A product of the simulated noise and the encoder output is then passed to the decoder.
-We use the counter decoding technique from the PT [38], to drive the decoder.
+We use the counter decoding technique from the PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden], to drive the decoder.
 The decoder learns to reconstruct the original face sequence and the input counter values.
-Thus, the model is trained with the following loss function;
+Thus, the model is trained end-to-end with the following loss function;
 $L_{Face}=\dfrac{1}{U_{f}}\sum_{u=1}^{U_{f}}((f_{u}-\hat{f_{u}})^{2}+\lambda_{CN}(c_{u}-\hat{c_{u}})^{2})$ (6)
@@
 Once fully trained we pass each codebook embedding, $F^{z}_{i}$, through the decoder to give the learnt dictionary of facial expressions in Euclidean space, $DF=[F_{1},F_{2},...,F_{N_{f}}]$.
-3.0.5 Duration Generation
-When ground truth timing information is not available we propose a novel sign segmentation approach based on the stitching method described above.
-Given the ground truth gloss labels, we generate the stitched sequence, $P_{stitch}$, but without step 4 (sign resampling).
-Comparing the stitch sequence and the ground truth, we find that the motion can vary due to different lexical variants present compared to our dictionary.
-However, we find that the handshape is often still consistent.
-So, we take the keypoints that correspond to the signer’s hands and normalise the rotation so that the index finger metacarpal bone is fixed on the y-axis and the palm is fixed on the xy-plane, giving $P^{H}_{stitch},P^{H}$.
-Our next step is to align the two sequences so that we can infer the duration of the signs in the ground truth.
-For this we apply Dynamic Time Warping (DTW), such that;
-$A_{i,j}=DTW(P^{H}_{stitch},P^{H})$ (7)
-As we know the duration of the isolated signs in the stitched sequence, by analysing the alignment path, $A_{j}$, we can infer the duration of the signs in the original ground truth sequence.
-3.0.6 Cutoff Generation
-Experiments reveal that each sequence contains a distinct range of frequencies correlated to the signer’s style.
-Fast motions contain high frequencies, while soft, slow signing involves lower frequencies, typically within the range of 1 to 25 Hz.
+3.5 Cutoff Generation
 To generate the ground truth cutoffs used in training, we once again apply our stitching approach.
@@
 We opt to use this method over just analyzing the frequency in the original sequence as we do not have an ideal filter.
-Thus, the butterworth filter has unintended effects on the frequency and phase blow the cutoff.
-4 Experimental Setup
-4.0.1 Implementation Details
-In our experiments, we conducted a grid search for optimal hyper-parameters and identified the following settings as the most effective.
-Our encoder-decoder translation model is constructed with an embedding size of 512 and a feed-forward size of 1024.
-We find that the optimal number of layers and heads is dataset-dependent, with smaller datasets requiring fewer layers compared to Meine DGS Annotated (mDGS) where we use 3 layers and 4 heads.
-The models utilise dropout with a probability of 0.1 [41], ReLU activations between layers [1], and pre-layer normalisation for regularisation and training stability.
-Training employs a ’reduce on plateau’ scheduler with a patience of 5 and a decrease factor of 0.8.
-The layers are initialised using a Xavier initializer [12] with zero bias, and during training, Adam optimization is employed [19].
-The initial learning rate is set to $10^{-4}$, and we train the model until convergence.
-During decoding, we utilise a greedy search algorithm.
-The loss scaling factors, $\lambda_{y}$, $\lambda_{d}$, $\lambda_{f}$ and $\lambda_{C}$ are set to 1.0, 0.1, 0.3 and 0.2, respectively.
-When stitching we enforce a minimum sign length, $U_{min}$ of 4 frames.
-The cropping threshold, $\alpha_{crop}$ is dataset-dependent, we find values between the range of 0.1 to 0.35 most effective.
-All sequences are subsampled to 12 frames per second (fps) for computational efficiency.
-For each dataset, we create a dictionary of 500 facial expressions.
-We scale the counter loss by 0.01 and we set an embedding dimension of 512.
-The encoder and decoder are initialized with the same settings as our translation model.
-The angular pose representation comprises 104 angles, while the Euclidean representation consists of 61 keypoints (21 for each hand and 19 for the body and face).
-The face mesh we add includes 128 keypoints, which are a subset of Mediapipe’s 478-face mesh [26].
-For comparison, we train a progressive transformer on each dataset until convergence using the parameters from [38].
-4.0.2 Datasets
-The approach is tested on three datasets, the Public Corpus of German Sign Language, 3rd release, the mDGS dataset [23], PHOENIX14T [5] and the BSL Corpus T (BSLCPT) [40].
-BSLCPT contains 211 participants from 8 regions in the UK, performing 4792 individual signs from a range of age groups.
-The participants perform narrative, interviews and participate in free conversation.
-Similarly, mDGS contains 330 participants engaging in free-form signing.
-Whereas, the PHOENIX14T dataset is extracted from German TV weather broadcasters and contains over 8,000 parallel sequences.
-4.0.3 Dictionary
-In the following experiment, we test two different dictionaries: 1) collected from isolated examples, and 2) a dictionary created from continuous data. Next, we provide further details about each:
-Isolated: Here, the signs are sourced from individuals who perform each sign in isolation, typically starting from and returning to a resting position.
-When experimenting on the BSLCPT we use the Signbank dataset [7], it contains over 3,000 signs and includes all the lexical variants found in the BSLCPT dataset.
-However no such dictionary exists for the PHOENIX14T and mDGS dataset, therefore we collect a dictionary from a range of sources such as [23].
-We find the mDGS dataset has a target vocabulary of 10,801.
-However, without the gloss variant, we find the core gloss vocabulary reduces to 4,434.
-We collect a total of 7,206 signs to experiment with.
-We use the method described in Sec.
-3 (Stitching), step 1 to overcome issues with an incomplete vocabulary.
-To create word embeddings we apply Facebook’s implementation of Fasttext [27].
-When experimenting on the BSLCPT we use the English implementation whereas on PHOENIX14T and mDGS we use the German implementation.
-Continuous: Here we create a dictionary using the gloss timing annotations.
-The signs are taken from the test and dev data only, so that the back translation model has not seen the signs during training.
-As the examples come from the continuous sign, we omit the cropping step of the stitching pipeline.
-These dictionaries have an abundance of signs to choose from when stitching.
-We filter the dictionary and remove short signs as these are most likely co-articulated and therefore not suitable or out of context.
-We opt to randomly select the first sign in the sequence, and subsequent signs are chosen to ensure the most natural transition.
-Therefore, we select the sign from the dictionary in which the location of the wrist is closest to the last frame of the previous sign.
-4.0.4 Evaluation Metrics
-To evaluate our approach we employ a CSLR model (Sign Language Transformers [6]) to conduct back-translation, the same as [14, 36, 39, 52].
-A model is trained for each dataset and the parameters are frozen so that results are consistent.
-BLEU [31], Rouge [25], and chrF [32] scores are computed between the predicted text and the ground truth.
-To evaluate the pose we employ Dynamic Time Warping Mean Joint Error (DTW-MJE), the metric compresses and stretches time to find the best alignment between the ground truth, $p$, and the predicted pose, $\hat{p}$.
+Thus, the Butterworth filter has unintended effects on frequencies below the cutoff.
 4 Experiments
+4.1 Implementation Details
+We apply the approach to three datasets, the Public Corpus of German Sign Language, 3rd release, the Meine DGS Annotated (mDGS) dataset [Konrad et al.(2020)Konrad, Hanke, Langer, Blanck, Bleicken, Hofmann, Jeziorski, König, König, Nishio, Regen, Salden, Wagner, Worseck, Böse, Jahn, and Schulder], RWTH-PHOENIX-Weather-2014T (PHOENIX14T) [Camgoz et al.(2018)Camgoz, Hadfield, Koller, Ney, and Bowden] and the BSL Corpus T (BSLCPT) [Schembri(2008)].
+To evaluate our approach we employ a CSLR model (Sign Language Transformers [Camgoz et al.(2020)Camgoz, Koller, Hadfield, and Bowden]) to conduct back-translation, the same as [Huang et al.(2021)Huang, Pan, Zhao, and Tian, Saunders et al.(2020a)Saunders, Camgoz, and Bowden, Saunders et al.(2021b)Saunders, Camgoz, and Bowden, Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu].
+BLEU [Papineni et al.(2002)Papineni, Roukos, Ward, and Zhu], Rouge [Lin(2004)], and chrF [Popović(2015)] scores are computed between the predicted text and the ground truth.
+Finally, to evaluate the pose we employ Dynamic Time Warping Mean Joint Error (DTW-MJE).
+In the following experiment, we test two different dictionaries: 1) collected from isolated examples, and 2) a dictionary created from continuous data.
+Noted as Isolated and continuous in the following tables.
+Further information about the datasets, dictionaries and model implementation can be found in the supplementary material.
 4.2 Quantitative Evaluation
-5.1.1 Text-to-Gloss Translation Results
-We start by evaluating the T2G translation performance described in Sec.
-3.0.1.
-Tab.
-1 shows the performance on all three datasets.
+Text-to-Gloss Translation Results: We start by evaluating the T2G translation performance described in Section 3.1.
+Table 1 shows the performance on all three datasets.
 We suggest that the difficulty of a dataset is proportional to the vocabulary and the total number of sequences used in training.
 We find the best performance on PHOENIX14T data which has the highest number of sequences per token, achieving 18.11 BLEU-4.
-In comparison to previous works, we find by having the model predict duration, face and cutoff we can achieve higher BLEU-1 scores, but at the cost of a lower BLEU-4 in comparison to [38].
+In comparison to previous works, we find by having the model predict duration, face and cutoff we can achieve higher BLEU-1 scores, but at the cost of a lower BLEU-4 in comparison to [Saunders et al.(2020c)Saunders, Camgoz, and Bowden].
 On the more challenging mDGS dataset we find a considerably lower BLUE-4 score due to the larger domain of discourse.
@@
 Overall we find the model to be performing as expected.
-Table 1: The results of translating from Text-to-Gloss on the BSL Corpus T, RWTH-PHOENIX-Weather-2014T and Meine DGS Annotated dataset.
 TEST SET DEV SET
@@
 PHOENIX14T 55.48 36.54 25.18 18.11 49.30 53.83 56.55 37.32 25.85 18.74 48.91 54.81
-PHOENIX14T [38] 55.18 37.10 26.24 19.10 - 54.55 55.65 38.21 27.36 20.23 - 55.41
-PHOENIX14T [42] 50.67 32.25 21.54 15.26 - 48.10 50.15 32.47 22.30 16.34 - 48.42
-5.1.2 Text-to-Pose Translation Results
-Note in the following experiments the back-translation model’s performance (shown as GT top row of Tab.
-2, 3 and 4) should be considered the upper limit of performance.
+PHOENIX14T [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 55.18 37.10 26.24 19.10 - 54.55 55.65 38.21 27.36 20.23 - 55.41
+PHOENIX14T [Stoll et al.(2018)Stoll, Camgöz, Hadfield, and Bowden] 50.67 32.25 21.54 15.26 - 48.10 50.15 32.47 22.30 16.34 - 48.42
+Table 1: The results of translating from Text-to-Gloss on the BSL Corpus T, RWTH-PHOENIX-Weather-2014T and Meine DGS Annotated dataset.
+Text-to-Pose Translation Results: Note in the following experiments the back-translation model’s performance (shown as GT top row of Table 2, 3 and 4) should be considered the upper limit of performance.
 In this section, we evaluate the Text-to-Pose (T2P) performance using back translation.
-To allow for a comparison we train two versions of the PT with the setting presented in [38].
+To allow for a comparison we train two versions of the PT with the setting presented in [Saunders et al.(2020c)Saunders, Camgoz, and Bowden].
 PT is the standard architecture, while PT + GN is trained with Gaussian Noise added to the input.
@@
 For back-translation, we outperform the PT on all metrics.
-Showing significant improvements in BLEU-1 score of 98% and 269% on the mDGS and BSLCPT dev set (comparing PT + GN and Stitcher (continuous), Tab.
-2 and 3).
-Deep learning models exhibit a bias toward the data they were trained on and often show poor out-of-domain performance.
-Unsurprisingly, the performance improves when using the continuous dictionary.
-We find only a small increase in BLEU-1 of 0.13 on the BSLCPT dev set (Tab.
-2), most likely due to the isolated dictionary containing the lexical variants found in the original data.
-Whereas we see a larger increase on the mDGS dataset (Tab.
-3).
-Table 2: The results of translating from Text-to-Pose on the BSL Corpus T dataset.
+Showing significant improvements in BLEU-1 score of 98% and 269% on the mDGS and BSLCPT dev set (comparing PT + GN and Stitcher (continuous), Table 2 and 3).
 BSLCPT TEST SET DEV SET
@@
 GT 0.000 17.3 3.96 1.37 0.54 13.00 21.76 0.000 17.32 3.71 1.08 0.39 13.04 21.89
-PT [38] 0.288 4.40 0.65 0.18 0.00 5.80 8.22 0.292 4.00 0.61 0.10 0.00 5.69 8.02
-PT + GN [38] 0.267 4.96 0.55 0.13 0.00 6.38 8.82 0.258 4.47 0.63 0.09 0.00 6.14 8.89
+PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.288 4.40 0.65 0.18 0.00 5.80 8.22 0.292 4.00 0.61 0.10 0.00 5.69 8.02
+PT + GN [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.267 4.96 0.55 0.13 0.00 6.38 8.82 0.258 4.47 0.63 0.09 0.00 6.14 8.89
 Stitcher (Isolated) 0.588 16.37 2.86 0.75 0.28 14.07 20.84 0.592 16.39 2.82 0.58 0.00 13.9 19.55
 Stitcher (continuous) 0.575 16.99 3.65 1.03 0.41 14.32 20.65 0.573 16.52 3.19 0.73 0.00 14.34 20.53
-Table 3: The results of translating from Text-to-Pose on the Meine DGS Annotated (mDGS) dataset.
+Table 2: The results of translating from Text-to-Pose on the BSL Corpus T dataset.
+Deep learning models exhibit a bias toward the data they were trained on and often show poor out-of-domain performance.
+Unsurprisingly, the performance improves when using the continuous dictionary.
+We find only a small increase in BLEU-1 of 0.13 on the BSLCPT dev set (Table 2), most likely due to the isolated dictionary containing the lexical variants found in the original data.
+Whereas we see a larger increase on the mDGS dataset (Table 3).
 mDGS TEST SET DEV SET
@@
 GT 0.000 20.87 5.60 1.89 0.80 17.56 23.78 0.000 20.75 5.43 1.81 0.76 17.63 23.41
-PT [38] 0.229 6.11 0.94 0.21 0.05 8.07 8.36 0.228 6.22 0.98 0.17 0.00 8.23 8.44
-PT + GN [38] 0.2245 7.18 1.48 0.40 0.01 8.46 8.38 0.2241 9.22 1.63 0.38 0.01 8.94 8.57
+PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.229 6.11 0.94 0.21 0.05 8.07 8.36 0.228 6.22 0.98 0.17 0.00 8.23 8.44
+PT + GN [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.2245 7.18 1.48 0.40 0.01 8.46 8.38 0.2241 9.22 1.63 0.38 0.01 8.94 8.57
 Stitcher (Isolated) 0.581 16.63 3.75 0.94 0.22 13.39 21.69 0.592 16.9 3.67 0.95 0.32 13.9 21.34
 Stitcher (Continuous) 0.637 18.64 4.17 1.07 0.39 16.86 21.80 0.637 18.27 4.07 1.19 0.43 16.75 21.25
-Previous work has primarily focused on G2P translation, therefore to facilitate a meaningful comparison we present two versions of the model.
-First, a G2P version, where we use the ground truth data and just apply the stitching module, and, secondly our T2P approach (translation then stitching).
-Results for comparison are provided by [52].
-We find our approach outperforms previous work on the BLEU-1 to 2 scores increasing the score by 56% and 19%, respectively (comparing Tab.
-4, row 7 and 11).
-We also find significant improvement in ROUGE and chrF metrics.
-Table 4: The results of translating from Gloss-to-Pose (G2P) and Text-to-Pose (T2P) on the RWTH-PHOENIX-Weather-2014T dataset.
+Table 3: The results of translating from Text-to-Pose on the Meine DGS Annotated (mDGS) dataset.
 PHOENIX14T
@@
 GT 0.000 32.41 20.19 14.41 11.32 33.84 32.96
-PT [38] 0.197 6.27 3.33 2.14 1.59 14.52 9.50
-PT + GN [38] 0.191 11.45 7.08 5.08 4.04 19.09 14.52
-NAT-AT [14] 0.177 14.26 9.93 7.11 5.53 21.87 18.72
-NAT-EA [14] 0.146 15.12 10.45 7.99 6.66 22.98 19.43
-PoseVQ-MP [52] 0.146 15.43 10.69 8.26 6.98 - -
-PoseVQ-DDM [52] 0.116 16.11 11.37 9.22 7.50 - -
+PT [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.197 6.27 3.33 2.14 1.59 14.52 9.50
+PT + GN [Saunders et al.(2020c)Saunders, Camgoz, and Bowden] 0.191 11.45 7.08 5.08 4.04 19.09 14.52
+NAT-AT [Huang et al.(2021)Huang, Pan, Zhao, and Tian] 0.177 14.26 9.93 7.11 5.53 21.87 18.72
+NAT-EA [Huang et al.(2021)Huang, Pan, Zhao, and Tian] 0.146 15.12 10.45 7.99 6.66 22.98 19.43
+PoseVQ-MP [Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu] 0.146 15.43 10.69 8.26 6.98 - -
+PoseVQ-DDM [Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu] 0.116 16.11 11.37 9.22 7.50 - -
 Stitching G2P (Isolated) 0.593 21.47 8.79 4.25 2.49 23.74 20.32
@@
 Stitching T2P (Continuous) 0.572 25.14 13.54 8.98 6.67 29.5 26.49
-Using a continuous dictionary we can outperform all models except for the VQ based approaches on BLEU-3 to 4.
-As the VQ model is learning sub-units of a gloss sequence we suggest this gives it an advantage on higher n-grams, as each token can represent multiple signs.
-5.1.3 Segmentation Results
-To evaluate the stitching segmentation approach we calculate the duration for each gloss in the mDGS dataset test set.
-We achieve a sign level frame F1-score of 0.6373 a similar score compared to [29] that achieves a top score of 0.63.
-Demonstrating the effectiveness of stitching for sign segmentation.
-It is worth noting our approach requires gloss information, but is computationally inexpensive compared to the LSTM used in [29].
+Table 4: The results of translating from Gloss-to-Pose (G2P) and Text-to-Pose (T2P) on the RWTH-PHOENIX-Weather-2014T test set.
+Previous work has primarily focused on G2P translation, therefore to facilitate a meaningful comparison we present two versions of the model.
+First, a G2P version, where we use the ground truth data and just apply the stitching module, and, secondly our T2P approach (translation then stitching).
+Results for comparison are provided by [Xie et al.(2022)Xie, Zhang, Li, Tang, Du, and Hu].
+Note that the previous approaches do not use a dictionary of signs and instead attempt to regress the pose directly from the spoken language.
+We find our approach outperforms previous work on the BLEU-1 to 2 scores increasing the score by 56% and 19%, respectively (comparing Table 4, row 7 and 11).
+We also find significant improvement in ROUGE and chrF metrics.
+Using a continuous dictionary we can outperform all models except for the Vector Quantisation (VQ) based approaches on BLEU-3 to 4.
+As the VQ model is learning sub-units of a gloss sequence we suggest this gives it an advantage on higher n-grams, as each token that is predicted can represent multiple signs.
 4.3 Qualitative Evaluation
-5.2.1 Visual Outputs
-To demonstrate the approach’s effectiveness, we present skeleton and video outputs for two sign languages (BSL and DGS) [^1] https://github.com/walsharry/Sign_Stitching_Demos.
-To enable a fair evaluation, we also share failure cases where incorrect predictions from the initial translation produce repetitive sequences.
+Visual Outputs: To demonstrate the approach’s effectiveness, we present skeleton and video outputs for two sign languages (BSL and DGS) [^1] https://github.com/walsharry/Sign_Stitching_Demos.
 Furthermore, in the supplementary material, we share visualizations of the produced sequence.
-5.2.2 Survey Results
-The survey presented a comparison to PT, followed by an ablation of different components of the stitching approach.
+Survey Results: The survey presented a comparison to PT, followed by an ablation of different components of the stitching approach.
 17% of people surveyed were native Deaf signers, while 34% were L2 signers or language learners.
 87.5% preferred our approach compared to the PT, while the rest selected no preference.
-100% of people agreed that applying the filter improved the realism compared to no filtering, while resampling was found to be less important, with 37.5% selecting no preference between the resampled and normal sequence.
-Interestingly, 41.7% preferred using an isolated dictionary, while 20.8% preferred the continuous dictionary and the rest indicated no preference.
+100% of people agreed that applying the filter improved the realism compared, while resampling was found to be less important, with 37.5% selecting no preference between the resampled and normal sequence.
 5 Conclusion
@@
 Finally, the user evaluation agrees with the quantitative results, indicating our approach can produce realistic expressive Sign language.
+6 Acknowledgement
+We thank Adam Munder, Mariam Rahmani, and Abolfazl Ravanshad from OmniBridge, an Intel Venture.
+We also thank Thomas Hanke and the University of Hamburg for the use of the Meine DGS Annotated (mDGS) data.
+This work was supported by Intel, the SNSF project ‘SMILE II’ (CRSII5 193686), the European Union’s Horizon2020 programme (‘EASIER’ grant agreement 101016982) and the Innosuisse IICT Flagship (PFFS-21-47).
+This work reflects only the author’s view and the Commission is not responsible for any use that may be made of the information it contains.